Worker A annually invests $1,000 in an IRA for nine years (ages 27 through 35) and never makes another contribution. Worker B annually invests $1,000 in an IRA for thirty years (ages 36 through 65). Which worker will have more in his or her account when he or she retires if they both earn 8 percent on their investments?
Answer #1
Worker A
The formula for the future value of an ordinary annuity, as opposed to an annuity due, is as follows:
P = PMT x ((((1 + r) ^ n) - 1) / r)
Where:
P = the future value of an annuity stream To be computed
PMT = the dollar amount of each annuity payment 1000
r = the effective interest rate (also known as the discount rate) 8%
n = the number of periods in which payments will be made 9
Amount at 35th birthday = PMT x ((((1 + r) ^ n) - 1) / r)
Amount at 35th birthday = 1000*((((1 + 8%) ^ 9) - 1) / 8%)
Amount at 35th birthday $ 12,487.56
Now this amount remains invested till his age is 65 so years 30
The amount at 65th birthday= Investment * (1+rate)^time
The amount at 65th birthday= 12487.56*(1+8%)^30
The amount at 65th birthday= $125,658.03
Worker B
P = the future value of an annuity stream To be computed
PMT = the dollar amount of each annuity payment 1000
r = the effective interest rate (also known as the discount rate) 8%
n = the number of periods in which payments will be made 30
Amount at 65th birthday = PMT x ((((1 + r) ^ n) - 1) / r)
Amount at 65th birthday = 1000*((((1 + 8%) ^ 30) - 1) / 8%)
Amount at 65th birthday $113,283.21
As we can see worker A will have more money than worker B.